An aptitude test has a mean of 220 and a standard deviation of 10, find the corresponding z score for: a) a test score of 232, b)a test score of 212.
Thanks much- I just am having an awful time with stats!

Answer by stanbon(48568)

An aptitude test has a mean of 220 and a standard deviation of 10,
find the corresponding z score for:
a) a test score of 232,
z(232) = (232-220)/10 = 12/10 = 1.2
------------------------------------------
b)a test score of 212.
z(212) = (212-220)/10 = -8/10 = -0.8
------------------------------------------
Thanks much- I just am having an awful time with stats!
----------------------
What seems to be the trouble? Are you taking the
course on line? At what level is the course being
offered? high school?, college? graduate?
Why do you have to take it? nursing?, business? other?
==========================================
Cheers,
Stan H.
----------------------------